Write the following in terms of natural units: i) 20 m·s−1 ii) 4×105 kg·m·s−1 iii) 6.626×10−34...

  1. Write the following in terms of natural units:
    1. i) 20 m·s1

ii) 4×105 kg·m·s1

iii) 6.626×1034 J·s
Answer #1

Natural units are physical units of measurements. For example, e is the natural unit for charge and c is the natural unit of speed.

1)

since the units are of speed hence we have to write it in terms of c

now 3 x 10^ 8 m/s = 1 c

hence 20 m/s = 20 * c / 3 x 10^8 = 6.67 x 10^-8 c

2)

4 x 10^5 kg*m/s

where h is reduced planck's constant , G is gravitational constant and c is speed of light

hence kgm/s = root (hc^3/G) = 6.528

hence 1kgm/s = 6.528

hence 4 x 10^5 kgm/s = 4 x 10^5 x 6.528 = 2.611 x 10^6

3)

6.626 x 10-34 Js

now 1 Js = kg*m^2/s = 1.05 x 10^-34 on planck's scale

hence 6.626 x 10^-34 units will be = 6.626 x 10^-34 x 1.05 x 10^-34 = 6.96 x 10^-68